OpenStudy (jamanorthi):

margaret has a bag containing 5 red chips and 5 blue chips. what is the probability of drawing two red chips in two draws without replacement??

OpenStudy (whpalmer4):

well, what's the probability of drawing a red chip on the first draw?

OpenStudy (jamanorthi):

1/2

OpenStudy (jamanorthi):

are u there

OpenStudy (whpalmer4):

yes, I am. Okay, after you've drawn a red chip, what is the probability of drawing a red chip going to be, if you don't put that red chip you drew back?

OpenStudy (jamanorthi):

4/9

OpenStudy (whpalmer4):

very good. so we have two dependent events, the first with P = 1/2 and the second with P = 4/9. How do we combine them?

OpenStudy (jamanorthi):

multiply them

OpenStudy (jamanorthi):

so its 2/9 right??

OpenStudy (whpalmer4):

Good. the result is?

OpenStudy (jamanorthi):

thanks you so much

OpenStudy (whpalmer4):

yes, 2/9.